OpenStudy (anonymous):

3/8=x/12 solve for x solve for x x+1/8=2x-5/2 3x/3x-36=-1/x solve for x 2x/16x+30=-1/x solve for x

OpenStudy (anonymous):

medals :)

OpenStudy (bahrom7893):

3/8 = x/12 multiply both sides by 12: 12*3/8 = x x = 4*3/2 x = 6

OpenStudy (anonymous):

Bahrom is on the case so ur good

OpenStudy (bahrom7893):

wait first one's wrong wth.. is wrong with me

OpenStudy (bahrom7893):

3/8 = x/12 x = 36/8 = 9/2

OpenStudy (anonymous):

dude wat u doing i had ur back

OpenStudy (bahrom7893):

rewrite the rest using the parenthesis please... i cant tell what the question is

OpenStudy (anonymous):

1 is x-9/2

OpenStudy (anonymous):

x=9/2 my bad

OpenStudy (anonymous):

solve for x (x+1/8)=(2x-5/2)

OpenStudy (bahrom7893):

not like that mariah... parenthesis so that i can tell the fractions apart, put the numerator in parenthesis..

OpenStudy (anonymous):

he means brackets, i think

OpenStudy (bahrom7893):

no i mean there's a difference between (x+1)/4 and x + 1/4

OpenStudy (anonymous):

oh, my bad

OpenStudy (anonymous):

yeh thats brackets

OpenStudy (anonymous):

Indeed!

OpenStudy (bahrom7893):

[] arent these brackets?

OpenStudy (anonymous):

well they are the same except these are square brackets or soming like that, plus why u askking me im a moron

OpenStudy (anonymous):

solve for x (x+1)/8=2x-5/2 (3x)/3x-36=(-1)/x solve for x (2x)/16x+30=-(1)/x solve for x

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

like that?

OpenStudy (bahrom7893):

no mariah...  would be written as (3x+y)/8